# Flyby-induced displacement: analytic solution

### v2

@@ Abstract
Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the hyperbolic Scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method applied to the Scarf potential. Pure displacement arises when the wave zone contains an integer number of half-waves. The results confirm the prediction of Zel’dovich and Polnarev.
Key words: gravitational waves, displacement memory effect, flyby, Nikiforov-Uvarov method
@@
preprint: arXiv: 2502.0136v2 [gr-qc]
I Analytic solution for flyby
Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 . Zel’dovich - Polnarev suggestion was confirmed numerically for the derived-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 . Their profile does not support an analytic solution, however it can be approximated by the derived Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,,$ (1)
which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 . This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves. In this note we present yet another approximation of the flyby profile, namely by the hyperbolic Scarf potential scarfpot ,
${\cal A}^{scarf}(U)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (2)
for which an analytic solution is found; it confirms the above conjecture.
For small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
@@ I Analytic solution for flyby
In detail, the geodesics are, in Brinkmann coordinates, determined by the a Sturm-Liouville-type transverse equations DM-2 ,
$\dfrac{d^{2}\!X^{i}}{dU^{2}}-\frac{1}{2}{\cal A}X^{i}=0\,,$ (3)
$i=1,2$, where $U$ behaves as non-relativistic time and ${\cal A}={\cal A}(U)$ is the profile of the gravitational wave DBKP ; DGH91 . These equations are completed with the DM boundary conditions,
@@
Below we show that the hyperbolic Scarf potential scarfpot ; Scrf1 , eqn. (2), allows us to find an analytic solution which confirms the quantisation conjecture above. Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 . The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ into a form which can be solved by Legendre polynomials DM-1 .
$(1+t^{2})\frac{d^{2}X}{dt^{2}}+t\frac{dX}{dt}-\frac{gt}{1+t^{2}}X=0\,,$ (5)
which is of the generalised hypergeometric type, and can be solved analytically by the Nikiforov-Uvarov method NikiUvar . Since a fairly detailed description of this simple, yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , we only provide a brief summary of the algorithm in the Appendix.
@@
For Eq. (5) in particular,
$\sigma=1+t^{2},\;\;\;\pi_{1}=t,\;\;\;\sigma_{1}=-gt\,.$ (7)
Then,
$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,,$ (8)
@@
$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8},$ (9)
and
$|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (10)
@@
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.$ (11)
@@
To have real solutions, we assume $k>0$, that is we choose upper sign in Eq.(9): $k=\frac{\sqrt{1+16g^{2}}-1}{8}.$ On the other side, the quantization condition Eq.(34) requires
@@
since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,.$ Therefore we get, after some algebra,
@@
and using Eq. (10) for $|g|/\sqrt{k}$, we get
$\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ (14)
and we see that we must choose the lower signs in Eq.(11) and Eq.(12) to get $\mp\epsilon=1$. Then the solution of Eq.(14) is
$k_{n}=n(n+1),$ (15)
and Eq.(10) determines the “magic” strengths of the Scarf potential Eq.(2), for which we have polynomial solutions for $y_{n}(t)$ labeled by the integer “quantum” number $n$,
$\;\;\displaystyle{|g_{n}|=(2n+1)\sqrt{n(n+1)}\;.}\;\;$ (16)
It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. Because of Eq. (15) and Eq. (16), we have,
@@
can be easily integrated and we get (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
@@
Therefore, the corresponding geodesic trajectory is given by
$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,.}\;\;$ (20)
The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (35) in which we need to find the weight function $\rho(t)$, for which we found,
Therefore
$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,.}\;\;$ (22)
The normalization constants $B_{n}$ should be determined from the initial conditions. The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$ They can be expressed in terms of Routh-Romanovski polynomials scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
For (half)wave numbers $n=1,2,3,4$ the analytic solutions (20) reproduce almost identically those half-numerical ones shown in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, DM requires to find a high number of decimals and the numerical procedure becomes tedious. Plotting instead the analytic solution (20)-(22) yields FIG.2.
Figure 2: DM geodesics for the Scarf profile (2) for wave numbers $n=5$ and $n=6$.
@@
In conclusion, our results show that the profile proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by the Scarf potential (2) in FIG.1 which supports analytic solutions, (20)-(22), labeled by an integer $n$, (16), which can be constructed by the Nikiforov-Uvarov method. The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.2, extending an observation made DM-2 for the Gaussian/Pöschl-Teller profile.
Further physical applications of the Scarf potential which include supersymmetry scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref. Raposo . Note that, in the supersymmetry literature, the Scarf hyperbolic potential has recently attracted much attention Scrf1 . This potential was discovered by Gendenshtein and is therefore referred also to as the Gendenstein potential Gendenon ; Natanson . See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for related work.
For isotropic profile we have a “universal model” : every plane gravitational wave can be mapped locally conformally to free Minkowski space by following Niederer, or Arnold Niederer ; Arnold ; Takagi , consistently with the vanishing of the Cotton or Weyl tensor Inomata ; Dhasmana ; GibbonsDark ; ZZH21 .
However gravitational waves are isotropic only in 1d, and their anisotropy in d=2 dim implies that we arrive to unchartered territory Lukash ; Inomata ; ZZH21 and it is not clear whether we still have a (conformal) “universal ancestor”.

### v3

# Flyby-induced displacement effect: an analytic solution

@@ Abstract
We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby. The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method. Our results confirm the prediction of Zel’dovich and Polnarev about certain “magical” amplitudes of the potential.
Key words: gravitational waves; displacement memory effect; flyby; Nikiforov-Uvarov method.
@@
preprint: arXiv: 2502.0136v3 [gr-qc]
I Introduction
One of the early proposals to observe gravitational waves , put forward by Braginsky and Thorne BraTho , was to study the motion of particles initially in rest after the passing of a burst of gravitational waves . They argued indeed that Burst With Memory could be among the earliest kinds of gravitational waves detected BraTho .
Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories. Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : …although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.
The GW memory effect offers a way to test general relativity including the nonlinear nature of gravity, and provides additional information about the GW source. Preliminary studies have suggested that the detection of GW memory from individual mergers of massive binary black holes is expected with LISA Favata . However, more realistic estimates show that observing the GW memory effect at detectors such as LISA, although feasible, is a challenging task fly-1 ; fly-2 . Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have predictably yielded negative results fly-3 ; fly-4 . However the prospects for the next generation of gravitational wave detectors are brighter even for future ground-based detectors such as Cosmic Explorer and the Einstein Telescope (see fly-2 ; fly-4 and references therein).
Zel’dovich and Polnarev did not present any proof. Although their suggestion could be confirmed in Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there were some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory . Then it came as a surprise that that the DM effect could be confirmed numerically DM-1 ; Jibril ; DM-2 for plane waves — provided the amplitude takes some “magical” values. Below we present an analytic proof.
In Brinkmann coordinates, the metric is written as Brinkmann ,
$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,.$ (1)
Gibbons and Hawking GibbHaw71 argued that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source GibbonsP . This lead them to propose the derivative of the Gaussian,
${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,.$ (2)
as profile.
The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,.$ (3)
Then the geodesics can be studied semi-analytically DM-1 ; DM-2 . Our investigations led us to conjecture that DM arises when a quantization holds, namely when the wave zone contains an integer number of what we called half-waves.
In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal. It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt1 .,
${\cal A}^{Scarf}(U)=2g\frac{\;\;d}{dU}\left(\frac{1}{\cosh U}\right)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (4)
to be compared with (3). For small values of $U$ the Scarf profile (4) behaves as the derivative of the Gaussian or of Pöschl-Teller , (2) or (3), but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1. Their similarity follows from that all three of them are derivatives of similar bell-shaped functions. We note also all three profiles in (2)-(3)-(4) are derivatives of an even function and are therefore odd.
@@ I Introduction
The geodesics are, in Brinkmann coordinates, determined by the transverse equations DM-2 ,
$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}X^{\pm}=0\,,$ (5)
where the $\pm$ label refers to the upper resp. lower component. The sign of the second  term comes from the relative minus between the coordinates in (1). The equation for the $V$ component follows from these DM-1 ; DM-2 , and is therefore omitted. We recall that $U$ behaves, in the Bargmann framework DBKP ; DGH91 , as non-relativistic time. These Sturm-Liouville-type equations are completed with the DM boundary conditions,
@@
Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above. Our clue is that Eqn.(5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its (non-normalizable) zero-energy ground state solutions DM-1 .
II Analytic solution for flyby
We first recall how things go for the simple Pöschl-Teller case. DM-1 ; DM-2 . Putting $t=\tanh U$, into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields
$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,,$ (7)
Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is
$g=-4n(n+1){\quad\text{\small where}\quad}n>0\quad\text{is a natural integer}\,.$ (8)
However passing to the other component, the linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$. But this is impossible because $g$ has already been fixed. In conclusion, one has what we called “half DM” DM-2 : we get a DM solution in one or in the other sector, but not for both in general — unless the “wrong” component is set to zero.
Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$), DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
$(1+t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}+t\frac{dX^{\pm}}{dt}{\mp}g\frac{t}{1+t^{2}}X^{\pm}=0\,.$ (9)
The two equations here are again uncoupled and can therefore be solved independently. Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term which compensates the $\pm$ sign change between the components. Therefore the same $g$ works for both components, as it will be seen in sec. III.
Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller . For derived Pöschl-Teller it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions in eqn. # (IV.16) of DM-2 .
Turning to Scarf, we note that (9) is of the generalised hypergeometric type which can be solved analytically by the Nikiforov-Uvarov (NU) method NikiUvar . Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main result. The algorithm, whose details are discussed in sec.III, yields the desired “magic” coefficient $g$ in (9),
$|g|=|g_{n}|=(2n+1)\sqrt{n(n+1)}\,,$ (10)
where $n$ is an integer. Then the trajectory is,
$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,}\;\;$ (11)
where $y_{n}(t)$ is the polynomial
$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,}\;\;$ (12)
with weight function
The normalization constants $B_{n}$ is determined from the initial conditions.
The analytic solutions (11) reproduce almost identically those obtained numerically in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, the numerical procedure becomes tedious as it requires to calculate a high number of decimals. For high wave numbers the result becomes also inaccurate because the procedure used in DM-2 is only an approximation. Plotting instead the analytic solution (11)-(12), FIG.3, is straightforward, underlining the advantage of the analytical solution.
Figure 2: The DM geodesics for the derived Pöschl-Teller (3) and for the Scarf profiles (4) are similar however not fully identical, as shown for wave numbers ${\bf n=1}$ and ${\bf n=2}$. (The trajectories were rescaled for better comparison).
Figure 3: DM geodesics for the Scarf profile (4) with wave numbers ${\bf n=5}$ and ${\bf n=6}$.
III Transverse geodesics by the Nikiforov-Uvarov method
@@
In the Scarp case Eq. (9) $\sigma=1+t^{2},\;\pi_{1}=t,\;\sigma_{1}=-gt\,.$ Then,
$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,.$ (15)
@@
$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8}{\quad\text{\small and}\quad}|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (16)
@@
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\quad$ (17)
@@
To have real solutions, we assume $k>0$, that is we choose th upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (36) requires
@@
since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,$ and we get, after some algebra,
@@
and using Eq. (16) for $|g|/\sqrt{k}$, we get $\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ and we see that we must choose the lower signs in Eq.(17) and Eq.(18) to get $\mp\epsilon=1$. Then the solution is $k_{n}=n(n+1)\,,$ and Eq.(16) determines the “magic” strengths labeled by the integer “quantum” number $n$, eqn. (10), for which we have polynomial solutions $y_{n}(t)$.
The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$. The first three of them are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
@@
can be easily integrated to yield (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
@@
Therefore, the corresponding geodesic trajectory is given by eqn. (11). The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (37) which requires to calculate the weight function $\rho(t)$ in (13). Then we end up with eqn. (12).
Eqn. (13) implies that $\rho(t)$ does not change when $t$ and $g$ change signs simultaneously, while due to the derivatives in eqn. (12), the polynomials $y_{n}(t)$ acquire a $(-1)^{n}$ factor. Then eqn. (11) implies the parity behaviour
$X^{\pm}(-t)=(-1)^{n}X^{\mp}(t)\,$ (23)
mentioned in sec.II. This curious pairing of the components is reminiscent of and is indeed closely related to supersymmetry MSUSY .
IV Conclusion
@@
The profile (2) proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4), as illustrated in FIG.1. The advantage of Scarf is that it supports analytic solutions, (11)-(12), labeled by an integer $n$, eqn (10). They are constructed by the Nikiforov-Uvarov method.
The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.3, extending the observation made in DM-2 for the Gaussian and Pöschl-Teller profiles. The origin of half-waves is clear: their number is that of the zeros of the polynomial $y_{n}(t)$.
Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 , as seen in FIG.2. It has also the advantage to be exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
The behavior of both the Scarf (4) and of the derived-Pöschl-Teller (3) is explained by their parity behaviour, (23), which explains in turn why do we get DM for both components. The same is true for any odd-order derivative of an even-profile function DM-1 ; DM-2 ; Benin . For even-order derivatives the analogous choices of the parameters would be contradictory, implying that one can gets DM only for one, but not for both components — unless the other is put to zero.
Physical applications of various incarnations of the Scarf potential are reviewed in ref. Raposo . They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt . The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY for gravitational waves with Scarf profile. See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.
We mention that our main tool here, namely the Nikiforov-Uvarov method, applies also to a number of other singular and periodic GW profiles, such as
${\cal A}=2g\left(1-\frac{1}{\tanh^{2}{U}}\right),\;\>{\cal A}=-2g\operatorname{cn}^{2}{\left(U,\frac{1}{\sqrt{2}}\right)},\;\;{\cal A}=2g\tan{U},\;\;{\cal A}=\frac{2g}{\tan{U}}\,,$ (24)
where $\operatorname{cn}(z,k)$ is Jacobian elliptic cosine function. Interestingly, the last profile corresponds to the Coulomb potential in constant curvature space $\mathbb{S}^{3}$, and hydrogenlike atoms in constant curvature spaces are most naturally considered by the Nikiforov-Uvarov method Alizzi .